OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 8x cos x - cos 8x sin x

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@LieutenantGeneral

hartnn (hartnn):

\(\Large \sin A \cos B - \cos A \sin B = \sin(A-B)\)

hartnn (hartnn):

A = 8x B =x

OpenStudy (lieutenantgeneral):

Sin (x+y) = sin(x)cos(y) + sin(y)cos(x) x = 8x y = 2x Sin (8x + 2x) = sin(10x) sin 8x cos 2x + cos 8x sin 2x = sin (8x + 2x) = sin 10x so the answer is A. identity: sin(A+B) = sinA cosB + cosA sinB

hartnn (hartnn):

but there is a negative sign in between!

OpenStudy (anonymous):

So it wouldnt be A? @hartnn

hartnn (hartnn):

A ???? you didn't list any choices

OpenStudy (anonymous):

sin 7x cos 7x cos 9x sin 9x

hartnn (hartnn):

and he just copied the answer from somewhere....

OpenStudy (anonymous):

oh..

hartnn (hartnn):

\(\Large \sin 8x \cos x - \cos 8x \sin x = \sin(8x-x) = ...\)

OpenStudy (anonymous):

sin7x

hartnn (hartnn):

correct! :)